Products > Test Equipment
Looking for the serive manual of an HP 70951 optical spectrum analyzer
<< < (6/8) > >>
_Wim_:

--- Quote from: nctnico on June 25, 2023, 07:49:26 pm ---Q403 could be a 2N7002 mosfet. It would be a good choice as a gain selector and a mosfet can be driven from a logic output directly (without a series resistor).

--- End quote ---

I think you are correct. I now start to doubt myself about the measured junction voltages, I will hopefully have some more time tomorrow evening. I will short the pads for low gain, this should then at least show a signal, albeit too low.
nctnico:
Removing Q403 would be a good idea. If it is broken, it could inject enough DC to overdrive the opamp to get it stuck to a rail.
_Wim_:

--- Quote from: nctnico on June 25, 2023, 09:26:59 pm ---Removing Q403 would be a good idea. If it is broken, it could inject enough DC to overdrive the opamp to get it stuck to a rail.

--- End quote ---

That was indeed what was happening. However U407 was also broken. After I replaced U407 and got the same result (railed to the positive rail) I wrongly assumed nothing was wrong with U407 and reinstalled the original part  :palm: Afterwards I removed Q403, this did not fix it also.

However both Q403 and U407 were broken! By removing Q403  and shorting Q403 to the lowest gain setting I now have a signal on lower sensitivity settings also. Off course some gain switching is not correct, I have some 2N7002 on order, as I did not have any suitable FET in stock.

But certainly good progress! Will off course also reinstall an AD797 for U407.

_Wim_:
A short update:
- Q403 has now been replaced by a 2N7002, this seems to work correctly, although I do not see much difference between the temporary short and with the 2N7002 in place (which is normal I guess, as it only adapts the gain slightly)
- broken analog output: is now fixed, this was also caused by a damaged AD797BR opamp (U105)
- non working transimpedance input: this was a user error, I did not configure the OSA correctly to use the external transimpedance input
- replaced U401 and U402 again by the original TL072, temperature went again above 100°C, which I guess is than normal (absolute max is 150°C according to the datasheet)

TODO's:
- verify all functionality
- find a keypad (not urgent as currently I am using the rubber buttons and a screenshot of the keypad)
_Wim_:
Some further updates that might be useful for others in the future:

- KE5FX 7470 plotter emulated works mostly correct to take screenshots (using an agilent 82357 GPIB usb adaptor)

Some information is the plotter screenshot is not correct or missing:
- Start,center and stop wavelengths are missing
- RBW is missing
- Video averaging number is correct, but is also shown when video averaging is off
- screenshot reference level is at top of display, while on the 70951 this is at the 9th graticule.

I have also done wavelength and power calibration using a 1550nm single mode laser. Wavelength was off by 0.6nm, power was off by 1.5dbm, but for power I am not really sure about the accuracy of my calibration. For wavelength I verified the result using an Acetylene reference cell (installed in the Anritsu MS9710C), and the result was bang on.



Navigation
Message Index
Next page
Previous page
There was an error while thanking
Thanking...

Go to full version
Powered by SMFPacks Advanced Attachments Uploader Mod